OpenStudy (anonymous):

1. The ordered pairs (1, 81), (2, 100), (3, 121), (4, 144), and (5, 168) represent a function. What is a rule that represents this function? (1 point) y = x^8 y = 8^x y = (x + 8)^2 y = 8x

OpenStudy (anonymous):

2. The ordered pairs (1, 6), (2, 36), (3, 216), (4, 1296), and (5, 7776) represent a function. What is a rule that represents this function? (1 point) y = 6x y = (x + 6)^2 y = x^6 y = 6^x 3. A snail travels at a rate of 2.35 feet per minute. Write a rule to describe the function. What is a rule that represents this function? (1 point) d = 5t; 11.74 ft. d = t + 2.35; 7.35 ft. d= t/2.35; 2.13 ft. d= 2.35t; 11.75 ft. Write a function rule for the area A of a triangle whose base, b, is 2 cm less than seven times the height, h. What is the area of the triangle when the height is 14 cm? (1 point) A = 7h^2 – 2h/2; 672 cm^2 A = 7h – 2/2; 48 cm^2 A = 7h – 2; 96 cm^2 A = 7h^2 = 2h; 1344 cm^2 5. Crystal earns $4.50 per hour mowing lawns. Write a rule to describe how the amount of money m earned is a function of the number of hours h spent mowing lawns. How much does Crystal earn is she works 1 hour and 45 minutes? (1 point) m = 4.50h; $7.88 m = h/ 4.50; $0.39 m = h + 45; $49.50 m = 4.50h; $6.53

OpenStudy (anonymous):

all you have to do is just plug values of x and y and check whether LHS is equal to RHS.. for all functions...